2x^2 - 9x -5=0
answer: x= -1/2 or 5

need a break down of how my text book got this answer

Answer by checkley77(12844)

x=-1/2 or 5
2x^2-9x-5=0
(2x+1)(x-5)=0
2x+1=0
2x=-1
x=-1/2 ans.
x-5=0
x=5 ans.
Proof:
2(-1/2)^2-9*-1/2-5=0
2*1/4+4.5-5=0
2/4+4.5-5=0
.5+4.5=5
5=5
2*5^2-9*5-5=0
2*25-45-5=0
50-50=0
0=0